Pls. help me solve these problem:
A grassy plot 25m by 30m is surrounded by a walk of uniform width. If the area of the walk is 300sq. m., how wide is it?

Answer by nerdybill(7384)

A grassy plot 25m by 30m is surrounded by a walk of uniform width. If the area of the walk is 300sq. m., how wide is it?
.
"area of grassy plot AND walk" minus "area of grassy plot" equals 300
.
Let x = width of walk
then
(25+2x)(30+2x) - (25)(30) = 300
(25+2x)(30+2x) - (25)(30) = 300
750+50x+60x+4x^2 - 750 = 300
50x+60x+4x^2 = 300
4x^2+110x = 300
4x^2+110x-300 = 0
2x^2+55x-150 = 0
2x^2+60x-5x-150 = 0
(2x^2+60x)-(5x+150) = 0
2x(x+30)-5(x+30) = 0
(2x-5)(x+30) = 0
x = {5/2, -30}
We can throw out the negative answer leaving:
x = 5/2 = 2.5 meters